OpenStudy (anonymous):

Will medal Part A: Write functions to represent Cory and Roger's collections throughout the years. (

OpenStudy (anonymous):

@dan815 @imqwerty @Owlcoffee @pooja195 @jhonyy9

OpenStudy (anonymous):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car.

OpenStudy (anonymous):

this goes with part A

OpenStudy (loser66):

Cory: at the beginning, he has 15 cars. After 1 year, he has 20% more. Now, 20% of 15 =??

OpenStudy (anonymous):

3?

OpenStudy (loser66):

Yes, so, after 1 year, he has 15 +3 = 18 cars, right?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

Now, year 2, at the beginning of year 2, he has 18 cars At the end of year 2, he has 20%of 18 =?

OpenStudy (anonymous):

so now how do i find rogers and then how would i put it into answer to the question

OpenStudy (anonymous):

3.6?

OpenStudy (anonymous):

is that 20% of 18

OpenStudy (loser66):

yes, so for the equation, you need something like \(C_n= 1.2C_{n-1}\) where \(C_0 = 15\)

OpenStudy (anonymous):

would c stand for cory and do you think it could be more explained, because that confused me a bit

OpenStudy (loser66):

C stands for the the cars Cory has. n stands for the year pass Like if n =1, that is year 1. The number of Cars he has \(C_1= 1.2 C_{1-1}=1.2C_0=1.2*15 =18\) if n=2, that is year 2, the number of Cars he has is \(C_2= 1.2 C_{2-1}=1.2 C_1=1.2*18 = 21.6\) As we do above 20% of 18 = 3.6 Hence after year 2, he has 18+3.6 =21.6

OpenStudy (loser66):

My equation is to generalize the formula used to calculate how many cars Cory has.

OpenStudy (anonymous):

so would those be the functions for cory

OpenStudy (loser66):

For Rogers, it is easy. At the beginning, he has 40 cars, each year he has 1 more. So \(C_n = 40 +n\) that is if n=1, after year 1, he has 40+1=41 after year 2, he has 40+2 =42 so far and so on.

OpenStudy (anonymous):

ok so now could you help me put those into an answer for my question .

OpenStudy (anonymous):

like if you looked at the question how would you answer it?

OpenStudy (loser66):

Honestly, I don't know what your prof wants you to do. To me, to answer those questions, I do like above.

OpenStudy (anonymous):

ok thx i will do my best then post the second part.